# Live magnetic observation of parahydrogen hyperpolarization dynamics

### v2

@@ Abstract
Hyperpolarized nuclear spins in molecules exhibit high magnetization that is unachievable by classical polarization techniques, making them widely used as sensors in physics, chemistry, and medicine. The state of a hyperpolarized material, however, is typically only studied indirectly and with partial destruction of magnetization, due to the nature of conventional detection by resonant-pickup nuclear magnetic resonance spectroscopy or imaging. Here we establish atomic magnetometers with sub- sensitivity as an use an alternative modality to detect in real time the complex dynamics of hyperpolarized materials without disturbing or interrupting the magnetogenesis process. As an example of dynamics that are impossible to detect in real time by conventional means, we examine parahydrogen-induced 1H and 13C magnetization during adiabatic eigenbasis transformations at -field avoided crossings. Continuous but nondestructive magnetometry reveals previously unseen spin dynamics, fidelity limits, and magnetization back-action effects. As a second example, we apply magnetometry to observe the chemical-exchange-driven 13C hyperpolarization of [1–13C]-pyruvate — the most important spin tracer for clinical metabolic imaging. Our approach can be readily combined with other high-sensitivity magnetometers and is applicable to a broader range of general observation scenarios involving production, transport and systems interaction of hyperpolarized compounds.
I Introduction
@@
Nuclear spin hyperpolarization is a family of non-equilibrium spin-manipulation methods, of interest across physics, chemistry, biomedicine and materials science. Hyperpolarization provides a route to extremely strong spin magnetization, boosting nuclear magnetic resonance (NMR) signals by orders of magnitude relative to equilibrium polarization by strong magnetic fields[1]. Hyperpolarized materials are used as particle-scattering targets[2, 3], clinical magnetic resonance imaging (MRI) probes of metabolism kinetics and metabolite biodistribution in vivo[4] and for characterizing structures of nanosheets, catalyst surface sites, and other low-abundance chemical samples[5]. Hyperpolarized materials also exhibit selective spin polarization, in which particular nuclei are strongly polarized while their neighbors remain unpolarized[6].
@@
Although powerful, capable of spectroscopically distinguishing nuclei by chemical shifts, conventional NMR imposes multiple limitations on the detection of hyperpolarized spins. Typically in NMR, one detects magnetization oscillations that are generated by tipping the magnetization off-axis by a radiofrequency (RF) pulse, then collecting the free-precession signal via RF inductive pickup. This “indirect” approach requires intervention for signal excitation, (i) imposing a time-frequency bandwidth limitation determined by the pulse-and-acquire repetition rate (see Figure 1), and (ii), consuming polarization in the process. It is common to use designed RF excitation pulse schemes to create or manipulate spin polarization[1, 7], but observing the intermediate dynamics with conventional NMR would require interrupting the pulse scheme for a detection period. The standard workaround is to repeat the experiment using multiple samples to create a time series, with each sample contributing one datapoint[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], but this comes at the cost of time, material, and the introduction of random error. Moreover such an approach cannot be used with one-time-hyperpolarized samples such as those destined for in vivo sensing[21], and those exhibiting complex or chaotic behavior[22].
@@
To overcome these limits and substantially widen the opportunities for sensing using hyperpolarized spins, an excitation-free detector is required, to sample the magnetization “directly”. In this work, we demonstrate an optically pumped magnetometer (OPM) as a suitable direct detector. As its sensing medium, the OPM employs the valence electrons of alkali-metal atoms (e.g., Rb) in the vapor phase[23]. Owing to the high magnetic moment, coherence time, and number of polarized spins in the vapor, OPMs can precisely measure the – fields produced even by thermally polarized NMR ensembles[24, 25, 26], over a range of frequencies from DC to kHz[27, 28, 10]. The OPM’s ability to non-destructively observe spin dynamics in real time opens a unique window into complex hyperpolarization processes. As we show, high-resolution magnetization tracking via OPMs can elucidate transfer mechanisms, allow speed optimization of transfer sequences, and potentially enable feedback control of hyperpolarization.
@@
These features are illustrated using parahydrogen-induced polarization (PHIP), a particularly powerful hyperpolarization technique for liquid samples that can polarize nuclear spins to nearly 100% in a range of systems, from dilute analytes to highly concentrated solutes[29]. PHIP operates by cooling H2 gas into a spin-singlet state, transferring the low-entropy spin order to the nuclei of a target molecule, and converting this spin order to vector-polarized form using control fields, which can range in strength from sub- [30, 31, 17, 18, 32, 16, 33] to T [34, 35, 36, 37]. Specifically, we use the OPM to demonstrate longitudinal magnetization detection during PHIP magnetogenesis: (1) nondestructively, (2) beyond the resolution-bandwidth limit of conventional NMR, where the sampling rate is much faster than the Larmor frequency, and (3) during ongoing adiabatic transitions. New features unveiled include the sensitivity of PHIP processes to spin-spin couplings, fidelity of repeated magnetizing and demagnetizing field cycles, and back-action effects of distant dipolar fields[38, 39, 40].
II Results
@@
In this section we highlight the unique features of direct sampling through the ability to unveil new information on magnetization dynamics in PHIP and related hyperpolarization processes. A more general interpretation of the capabilities follows in the Discussion (section III).
II.1 Observing Magnetogenesis during Adiabatic Eigenbasis Transformations
@@
Magnetometry and field cycling involves remarkably simple hardware comprising the OPM, a solenoid coil, a low-voltage programmable waveform generator and a passive magnetic shield, (see Figure 2b and Methods section). From a geometric standpoint, the solenoid coil is highly accommodating towards integrating the OPM; in an ideal scenario, the enclosed sample encounters a uniform magnetic field due to the solenoid, while the external OPM experiences no field.[42] Consequently, the impact of the field cycle on the magnetometer’s dynamic range is zero. Real-world conditions deviate only slightly from this ideal; the stray field generated by the solenoid is at least 500 times weaker than the internal field, so the OPM response is linear across several tens of in terms of the background field at the hyperpolarized sample.
To measure the field produced by the sample without background field contributions or other slow drifts in the OPM, the sign of the nuclear magnetization is repeatedly toggled via a series of high-fidelity pulses (see Figure 2c). When toggling is applied before the field cycle, no change is seen in the overall signal – confirming the initial hyperpolarized state is nonmagnetic. In contrast, after magnetization, a clear alternating pattern is seen where the peak-to-peak amplitude () equals twice the field produced by the sample. The 13C polarization for these experiments was measured separately in a high-field NMR system to be approximately 40%. Spin-selective toggle pulses allow magnetization from different nuclei to be selected, such as in Figure 2c where the pulses are selective to 1H. This reveals that the observed magnetic field is weighted more towards the 1H magnetization – as expected from the higher gyromagnetic ratio of 1H – even though the field sweep should polarize 1H and 13C equally (but in opposite directions). A small 13C magnetization is evident from the coherent oscillations superposed upon the toggled 1H signal, at the 13C Larmor frequency .
@@
Direct magnetization tracking immediately reveals that the speed of magnetogenesis in the linear-ramp case is limited by adiabaticity requirements, particularly around the level anticrossing region. Enhanced or ‘constant adiabaticity’ ramp profiles can be calculated to pass quickly through regions of the ramp where the eigenbasis does not change with field, but slowly through the important anticrossing region[43]. The OPM signal recorded during such a profile (see Figure 3a) shows a more uniform magnetization build-up rate compared to the linear profile. The increased average speed also helps to reduce relaxation losses, though these are not evident for fumarate because the ramp duration is much shorter than the 1H magnetization decay time constant , see Figure 5.
For the constant-adiabaticity case, our approach reveals expected[44] but previously unseen oscillations in the magnetization build-up curve, which appear to have a near-constant frequency across most of the sweep profile. The slow ones indicate non-adiabatic behavior and are known as “cycloid arcs” when the state of the 2-level system is represented as a trajectory on a Bloch sphere (see Figure 3d). Here, the cycloid arcs are mainly due to small coherences between states and that are introduced during the initial field switch to . The faster oscillations are quantum beats between states of the system with equal -projection angular momentum ( and ) that would not be present in the ideal polarization process. These and other features are visible even at low concentrations of the hyperpolarized product, including natural 13C isotopic abundance, due to the low total noise ( pTRMS, 0–100 Hz) of the OPM. The exact shape of the peak/trough features can vary with physical conditions such as pH and temperature, which can alter, e.g., the asymmetry of the heteronuclear -couplings. Real-time observation can be used to understand these dependencies, but also to minimize their nuisance effects: the magnetization process could be stopped at a beat’s peak, for example, thus optimizing the polarization yield at any pH and temperature. In complex, multi-spin systems, these features could also be used to understand magnetogenesis behavior that simplistic but more intuitive models (e.g., only involving a few spins) do not account for.
@@
II.2 Fidelity of Demagnetization and Remagnetization
@@
II.3 Quantifying the Sample’s Internal Magnetic Field and Backaction on Magnetogenesis
Nowadays, just as in the beginnings of pulsed NMR where RF radiation-damping effects were being noticed in concentrated liquids[38, 39], (e.g., water), practitioners of hyperpolarization are quickly realizing the importance of a sample’s own magnetic moment or “distant dipolar field” (DDF) that behaves as an additional temporal, spatially nonuniform and possible chaotic[45] magnetic field experienced by the spins. The DDF may interfere with the magnetization process itself[40] when comparable in size to the applied magnetic fields. Evidence for these effects can be observed for high (100 mM) concentrations of the hyperpolarized [1–13C]-fumarate isotopolog as shown in Figure 4, which compares the magnetization trajectories between samples of different labeling (2%, 20% and 99% [1-13C]). For dilute samples (e.g., 2% and 20% [1–13C]), the experimental data and simulated trajectories agree well, but beyond this (e.g., 99% [1–13C]), the number of magnetized molecules is sufficiently high for dipolar-field effects to be non-negligible. The build-up curves are noticeably smoother, with the cycloid arc features being less pronounced. This is attributed to destructive interference caused by the inhomogeneity of the DDF, and can be confirmed by comparing with numerically simulated profiles (inhomogeneity around ). This interference has been shown in other work to strongly limit polarization in PHIP experiments[40], but could only be studied after – not during – magnetogenesis. Direct tracking is advantageous because it allows one to understand and quantify DDF effects in real-time.
@@ II.3 Quantifying the Sample’s Internal Magnetic Field and Backaction on Magnetogenesis
II.4 Observing Magnetogenesis during Chemical Exchange at Fixed Field
@@
Polarization may again be probed in situ in a more conventional style[8, 9, 51, 12, 13] by interrupting the sub--field polarization process and carrying out a Fourier-transform NMR experiment to quantify the amplitude of spin-specific resonances. Spectra (see Figure 6b, recorded by the OPM at a solenoid field provide high-resolution information on the heteronuclear-coupled spin states, revealing in-phase and anti-phase spin multiplets around the 13C and 1H Larmor frequencies, respectively.
@@ III Discussion and Conclusion
We expect that the information provided by real-time observation techniques can facilitate new attempts to polarize specific molecules. A candidate with high potential as a multi-pathway reporter of metabolic disease by MRI is 13C-labeled glucose[54]. Unlike pyruvate, this is significantly far from preclinical and clinical applications, in part due to the large number of spins that result in complex conditions for efficient hyperpolarization.
@@
Besides use as a general method to study, machine-learn or optimize hyperpolarization, we expect the information provided by our approach to stimulate new ways to exploit hyperpolarized substances as sensors. For instance, the reversible field-sweeping experiment of Figure 4 shows [1–13C]-fumarate is a source of strong 1H/13C magnetization that can be switched on and off repeatedly. We speculate that this system could offer new opportunities for co-magnetometry, or be envisioned as a togglable polarized-nuclear target that is also lightweight and mobile enough to be used where current technologies do not reach, e.g., in space.

### v3

@@ Abstract
Significance statement: Molecules can be prepared in a hyperpolarized state to enhance their nuclear magnetic resonance (NMR) signals by four to five orders of magnitude, enabling many applications in spectroscopy and imaging (MRI).
But signal read-out typically requires a magnetic field pulse to perturb the system’s quantum state, followed by signal detection.
This process destroys the hyperpolarized state, and provides low-frame-rate observations when tracking dynamic processes.
Here we use optical magnetometry to observe induced quantum state transformations of hyperpolarized molecules in real time. This method offers an unprecedented view of hyperpolarization processes, revealing details analogous to switching from still-image film to video. Observation throughout the hyperpolarization process, even in the presence of time-dependent fields, opens possibilities for real-time control.
@@
Abstract: Hyperpolarized nuclear spins in molecules exhibit high magnetization that is unachievable by classical polarization techniques, making them widely used as sensors in physics, chemistry, and medicine. The state of a hyperpolarized material, however, is typically only studied indirectly and with partial destruction of magnetization, due to the nature of conventional detection by resonant-pickup nuclear magnetic resonance spectroscopy or imaging. Here we establish atomic magnetometers with sub- sensitivity as an alternative modality to detect in real time the complex dynamics of hyperpolarized materials without disturbing or interrupting the magnetogenesis process. As an example of dynamics that are impossible to detect in real time by conventional means, we examine parahydrogen-induced 1H and 13C magnetization during adiabatic eigenbasis transformations at -field avoided crossings. Continuous but nondestructive magnetometry reveals previously unseen spin dynamics, fidelity limits, and magnetization backaction effects. As a second example, we apply magnetometry to observe the chemical-exchange-driven 13C hyperpolarization of [1–13C]-pyruvate — the most important spin tracer for clinical metabolic imaging. The approach can be readily combined with other high-sensitivity magnetometers and is applicable to a broader range of general observation scenarios involving production, transport and systems interaction of hyperpolarized compounds.
@@
Nuclear spin hyperpolarization is a family of non-equilibrium spin-manipulation methods, of interest across physics, chemistry, biomedicine and materials science. Hyperpolarization provides a route to extremely strong spin magnetization, boosting NMR signals by orders of magnitude relative to equilibrium polarization by strong magnetic fields[1]. Hyperpolarized materials are used as particle-scattering targets[2, 3], clinical MRI probes of metabolism kinetics and metabolite biodistribution in vivo[4] and for characterizing structures of nanosheets, catalyst surface sites, and other low-abundance chemical samples[5]. Hyperpolarized materials also exhibit selective spin polarization, in which particular nuclei are strongly polarized while their neighbors remain unpolarized[6].
@@
Although powerful, capable of spectroscopically distinguishing nuclei by chemical shifts, conventional NMR imposes multiple limitations on the detection of hyperpolarized spins. Typically in NMR, one detects magnetization oscillations that are generated by tipping the magnetization off-axis by a radiofrequency (RF) pulse, then collecting the free-precession signal via RF inductive pickup. This “indirect” approach requires intervention for signal excitation, (i) imposing a time-frequency bandwidth limitation determined by the pulse-and-acquire repetition rate (Figure 1), and (ii), consuming polarization in the process. It is common to use designed RF excitation pulse schemes to create or manipulate spin polarization[1, 7], but observing the intermediate dynamics with conventional NMR would require interrupting the pulse scheme for a detection period. The standard workaround is to repeat the experiment using multiple samples to create a time series, with each sample contributing one datapoint[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], but this comes at the cost of time, material, and the introduction of random error. Moreover such an approach cannot be used with one-time-hyperpolarized samples such as those destined for in vivo sensing[22], and those exhibiting complex or chaotic behavior[23].
@@
To overcome these limits and substantially widen the opportunities for sensing using hyperpolarized spins, an excitation-free detector is required, to sample the magnetization directly. In this work, we demonstrate that an optically pumped magnetometer (OPM) as a suitable direct detector. As its sensing medium, the OPM employs the valence electrons of alkali-metal atoms (e.g., Rb) in the vapor phase[24]. Owing to the high magnetic moment, coherence time, and number of polarized spins in the vapor, OPMs can precisely measure the fT-pT fields produced by hyperpolarized[25, 8, 9, 26, 27, 28, 14] or thermally-polarized NMR ensembles[29, 30, 31], over a range of frequencies from DC to kHz[32, 33, 10]. The OPM’s ability to non-destructively observe spin dynamics in real time opens a unique window into complex hyperpolarization processes. As we show, high-resolution magnetization tracking via OPMs can elucidate transfer mechanisms, allow speed optimization of sequences, and potentially enable feedback control of hyperpolarization.
@@
These features are illustrated using parahydrogen-induced polarization (PHIP),[34, 35, 36] a particularly powerful hyperpolarization technique for liquid samples that can polarize nuclear spins to nearly 100% in a range of systems, from dilute analytes to highly concentrated solutes[37]. PHIP operates by cooling H2 gas into a spin-singlet state, transferring the low-entropy spin order to the nuclei of a target molecule, and converting this spin order to vector-polarized form using control fields, which can range in strength from sub- [38, 39, 17, 18, 40, 16, 41] to T [42, 43, 44, 45]. Specifically, we use the OPM to demonstrate longitudinal magnetization detection during PHIP magnetogenesis: (1) nondestructively, (2) beyond the resolution-bandwidth limit of conventional NMR, where the sampling rate is much faster than the Larmor frequency, and (3) during ongoing adiabatic transitions. Features unveiled include the sensitivity of PHIP processes to spin-spin couplings, fidelity of repeated magnetizing and demagnetizing field cycles, and backaction effects of distant dipolar fields[46, 47, 48].
Results
@@
In this section we highlight the unique features of direct sampling through the ability to unveil information on magnetization dynamics in PHIP and related hyperpolarization processes. A more general interpretation of the capabilities follows in the Discussion (Discussion and Conclusion).
.1 Observing Magnetogenesis during Adiabatic Eigenbasis Transformations
@@
Magnetometry and field cycling involves remarkably simple hardware comprising the OPM, a solenoid coil, a low-voltage programmable waveform generator and a passive magnetic shield, (Figure 2b and Methods section). From a geometric standpoint, the solenoid coil is highly accommodating towards integrating the OPM; in an ideal scenario, the enclosed sample encounters a uniform magnetic field due to the solenoid, while the external OPM experiences no field.[50] Consequently, the impact of the field cycle on the magnetometer’s dynamic range is zero. Real-world conditions deviate only slightly from this ideal; the stray field generated by the solenoid is at least 500 times weaker than the internal field, so the OPM response is linear across several tens of in terms of the background field at the hyperpolarized sample.
To measure the field produced by the sample without background field contributions or other slow drifts in the OPM, the sign of the nuclear magnetization is repeatedly toggled via a series of high-fidelity composite pulses (Figure 2c). When toggling is applied before the field cycle, no change is seen in the overall signal – confirming the initial hyperpolarized state is nonmagnetic. In contrast, after magnetization, a clear alternating pattern is seen where the peak-to-peak amplitude () equals twice the field produced by the sample. The 13C polarization for these experiments was measured separately in a high-field NMR system to be approximately 40%. Spin-selective toggle pulses allow magnetization from different nuclei to be inverted, such as in Figure 2c where the 90X180Y90X composite pulses are applied to give a net rotation of on 1H; these are mostly selective to 1H magnetization, since they yield a 22.5X45Y22.5X or around net rotation of the 13C magnetization, which quickly loses coherence after a few toggling repetitions. This reveals that the observed magnetic field is strongly weighted towards the 1H magnetization – as expected from the higher gyromagnetic ratio of 1H – even though the field sweep should polarize 1H and 13C equally (but in opposite directions). A small 13C magnetization is evident from the coherent oscillations superposed upon the toggled 1H signal, at the 13C Larmor frequency .
@@
Direct magnetization tracking immediately reveals that the speed of magnetogenesis in the linear-ramp case is limited by adiabaticity requirements, particularly around the level anticrossing region. Enhanced or ‘constant adiabaticity’ ramp profiles can be calculated to pass quickly through regions of the ramp where the eigenbasis does not change with field, but slowly through the important anticrossing region[51]. The OPM signal recorded during such a profile (see Figure 3b) shows a more uniform magnetization build-up rate compared to the linear profile. The increased average speed also helps to reduce relaxation losses, though these are not evident for fumarate because the ramp duration is much shorter than the 1H magnetization decay time constant , see Figure 5.
For the constant-adiabaticity case, our approach reveals expected[52] but previously unseen oscillations in the magnetization build-up curve, which appear to have a near-constant frequency across most of the sweep profile. The slow ones indicate non-adiabatic behavior and are known as “cycloid arcs” when the state of the 2-level system is represented as a trajectory on a Bloch sphere (see Figure 3d). Here, the cycloid arcs are mainly due to small coherences between states and that are introduced during the initial field switch to . The faster oscillations correspond to a weakly-populated magnetic state that precesses around the applied field (which for the majority of the sweep is 400 nT). This state is formed by the toggling pulses prior to the field sweep, which convert a small fraction of the population imbalance between 1H and states into coherences. Simulations show that these oscillations are predominantly sensed in the and axes, and are smaller if the toggling pulses are omitted. These and other features are visible even at low concentrations of the hyperpolarized product, including natural 13C isotopic abundance, due to the low total noise ( pTRMS, 0–100 Hz) of the OPM. The exact shape of the peak/trough features can vary with physical conditions such as pH and temperature, which can alter, e.g., the asymmetry of the heteronuclear -couplings. Real-time observation can be used to understand these dependencies, but also to minimize their nuisance effects: the magnetization process could be stopped at a beat’s peak, for example, thus optimizing the polarization yield at any pH and temperature. In complex, multi-spin systems, these features could also be used to understand magnetogenesis behavior that simplistic but more intuitive models (involving only a few spins) do not account for.
@@
.2 Fidelity of Demagnetization and Remagnetization
@@
A clear contrast can be seen between the magnetization profiles for sweeps of 0.5 vs. 2 s duration. The former appear smooth, whereas the latter show faster () oscillations on top, which are cycloid arcs due to coherences between states and . The 1 s sweep profile shows cycloid arc oscillations that appear to amplify over time, and eventually dominate the magnetization profile. This interesting behaviour is predicted by simulations, and is due to the sweep repetition rate coinciding with an overtone of the cycloid arc frequency, producing a resonance match. Simulations show that this effect may be suppressed by changing the field to avoid resonance. This can be verified both experimentally and by simulation; the inset in Figure 4 shows a smoother magnetization trajectory when the field during the pause is changed from to .
.3 Quantifying the Sample’s Internal Magnetic Field and Backaction on Magnetogenesis
Nowadays, just as in the beginnings of pulsed NMR where RF radiation-damping effects were being noticed in concentrated liquids[46, 47], (e.g., water), practitioners of hyperpolarization are quickly realizing the importance of a sample’s own magnetic moment or “distant dipolar field” (DDF) that behaves as an additional temporal, spatially nonuniform and possible chaotic[53] magnetic field experienced by the spins. The DDF may interfere with the magnetization process itself[48] when comparable in size to the applied magnetic fields. Evidence for these effects can be observed for high (100 mM) concentrations of the hyperpolarized [1–13C]-fumarate isotopolog as shown in Figure 4, which compares the magnetization trajectories between samples of different labeling (2.2%, 20% and 99% [1-13C]). For dilute samples (e.g., 2.2% and 20% [1–13C]), the experimental data and simulated trajectories agree well, but beyond this (e.g., 99% [1–13C]), the number of magnetized molecules is sufficiently high for dipolar-field effects to be non-negligible. The build-up curves are noticeably smoother, with the cycloid arc features being less pronounced. This is attributed to destructive interference caused by the inhomogeneity of the DDF, and can be confirmed by comparing with numerically simulated profiles (inhomogeneity around ). This interference has been shown in other work to strongly limit polarization in PHIP experiments[48], but could only be studied after – not during – magnetogenesis. Direct tracking is advantageous because it allows one to understand and quantify DDF effects in real-time, which is subject to investigation in ongoing studies.
@@
.4 Observing Magnetogenesis during Chemical Exchange at Fixed Field
@@
Polarization may again be probed in situ in a more conventional style[8, 9, 59, 12, 13] by interrupting the sub--field polarization process and carrying out a Fourier-transform NMR experiment to quantify the amplitude of spin-specific resonances. Spectra (Figure 6b) recorded by the OPM at a solenoid field provide high-resolution information on the heteronuclear-coupled spin states, revealing in-phase and anti-phase spin multiplets around the 13C and 1H Larmor frequencies, respectively.
@@
We expect that the information provided by real-time observation techniques can facilitate fresh attempts to polarize specific molecules. A candidate with high potential as a multi-pathway reporter of metabolic disease by MRI is 13C-labeled glucose[62]. Unlike pyruvate, this is significantly far from preclinical and clinical applications, in part due to the large number of spins that result in complex conditions for efficient hyperpolarization.
@@
Besides use as a general method to study, machine-learn or optimize hyperpolarization, we expect the information provided by our approach to stimulate original ways to exploit hyperpolarized substances as sensors. For instance, the reversible field-sweeping experiment of Figure 4 shows [1–13C]-fumarate is a source of strong 1H/13C magnetization that can be switched on and off repeatedly. We speculate that this system could offer new opportunities for co-magnetometry (with simultaneous magnetic field sensing using the electrons in the OPM and the nuclear spins in the sample, or the co-located 1H and 13C nuclear spin species, which are strongly or weakly coupled depending on the magnetic field strength), or be envisioned as a toggleable polarized-nuclear target that is also lightweight and mobile enough to be used where current technologies do not reach, e.g., in space.
Materials and Methods
Parahydrogen-induced polarization:
Hydrogen gas (H2) was prepared at 99% para enrichment using a commercial generator (ARS Cryo) operating at 30 K by flowing the gas over a hydrated iron (III) oxide catalyst (Merck KGaA, CAS 20344-49-4). The gas was stored at 10 bar in an aluminium cylinder and used throughout the day for experiments.
PHIP polarization of [1–13C]-fumarate and [1–13C]-maleate: To produce hyperpolarized [1–13C]-fumarate, the precursor solution for all experiments was 125 mM [1–13C]-acetylenedicarboxylic acid disodium salt (Merck), 125 mM sodium sulfite, and 3.5 mM ruthenium catalyst (pentamethyl cyclopentadienyl tris(acetonitrile) ruthenium(II) hexafluorophosphate) (Merck, CAS 99604-67-8) in 99% D2O, at pH 10-11. For each measurement, 600 µL of the solution was loaded by pipette into a 5 mm screw-top NMR tube, and the tube was sealed. The screw top had two 1/16" (o.d.) PTFE capillaries passing through, acting as an inlet and an outlet for the hydrogen gas. The tube was pressurized to 5 bar with parahydrogen, and its lower end placed into a mineral oil bath for the solution to warm (, 10 s). While still in the hot bath, parahydrogen was flowed through the tubes and bubbled through the solution (5 bar for 60 s). In most cases, the solution turned a pink/red color, indicating the reaction was near completion. The tube was depressurized, opened, and the solution was extracted through a plastic capillary into a 1 mL syringe. The solution was then injected by hand through a 1/16" (o.d.) PTFE capillary tube into a 2 mL, 8-425 vial (Merck) located inside the shielded apparatus of Figure 2b; the approximate delay between the end of the reaction and the start of field sweeps was 30 s.
To produce hyperpolarized [1–13C]-maleate, a similar procedure was used. The precursor solution was 125 mM [1–13C]-acetylene dicarboxylic acid and 5 mM rhodium catalyst ([1,4-bis(diphenylphosphino) butane] (1,5-cyclooctadiene) rhodium(I) tetrafluoroborate) (Merck, CAS
79255-71-3) in acetone-d6. 600 µL of this solution was loaded into a 5 mm screw-top NMR tube, and the same bubbling system was used for the reaction. Instead of using an oil bath, the tube was pressurized to 5 bar and then heated for 10 s with a heat gun until the solvent was boiling. Parahydrogen was then flowed through the sample for 45 s under continual heating. After the reaction, the tube was cooled in water (2 s) and then depressurized and opened to extract the sample in the same manner as for fumarate. The sample was also injected into the magnetometer apparatus as described before.
SABRE-SHEATH polarization of [1–13C]-pyruvate:
To hyperpolarize [1–13C]-pyruvate using SABRE-SHEATH, a solution of 60 mM [1–13C] sodium pyruvate (Merck, CAS 87976-71-4), 6 mM IrIMes(COD)Cl and 40 mM dimethyl sulfoxide-d6 in methanol-d4 was prepared. 1 mL of this sample was loaded into a 2 mL sample vial, the vial was sealed with a screw cap, and then placed into the shielded solenoid adjacent to the magnetometer. The sample vial cap featured a coaxial capillary (1/16" O.D.) with an inner glass capillary (0.15 mm o.d.) extending into the vial. These were used to bubble parahydrogen through the sample at 5 bar. The catalyst was activated by 10 mins of H2 bubbling and the sample was cooled to approximately 10 ºC by flowing ice-water around the sample vial. The sample was used under these conditions for the reported experiments.
Field control:
For magnetic field control and shielding, a multilayer MuMetal magnetic shield in a cylinder shape was used (Twinleaf model MS-1F, dimensions 30 cm length and 25 cm diameter). Inside the shield, a solenoid coil (7.5 mT/A, 15 mm diameter) was positioned parallel to the cylinder axis. Ramped magnetic fields were generated using a 12-bit digital-to-analog converter (DAC) integrated circuit (MCP4822, MicroChip Technology Inc.) whose output was connected to the solenoid. The DAC was digitally interfaced with a microcontroller, provided as part of the “NMRduino” open-source magnetic resonance platform[66]. The linear and constant-adiabaticity ramp profiles each comprised 800 equally spaced voltage points on a 1--precise time base. Additional magnetic coils located inside the shield were used to adjust the background field at the magnetometer, with bias currents zeroed before the experiments, controlled by a low-noise controller (Twinleaf CSB-10).
Optical magnetometry: The OPM used for detection comprised a cube-shaped alkali-metal vapor cell as its sensitive element ( mm3), inside which 87Rb and N2 buffer gas was contained and heated to . A D1-resonant (), circularly polarized light beam was passed through one of the cube faces to optically pump the atomic spin polarization of the vapor. A second, off-resonance beam was passed through one of the other faces, and out of the opposing side, to probe the atomic polarization via optical rotation.
The remaining cube face was placed against the exterior of the solenoid coil, near to the NMR sample vial contained within. A water cooling system was used to maintain a solenoid and vial temperature below (or in the case of the [1–13C]-pyruvate experiments), which resulted in a standoff distance of 5-6 mm between the vapor cell and the vial. Further details of the magnetometer can be found in past work[31].
Data acquisition and processing:
The rotation angle of the OPM probe beam was detected at a balanced polarimeter (Thorlabs PDB210A), which produced a differential output voltage linearly proportional to the - and -axis magnetic field amplitude, as illustrated in Figure 2b. The voltage was digitized by a second “NMRduino”[66] device (16 bit, , 1–5 kHz sampling rate) for a few seconds to minutes, depending on the experiment. The signal was streamed to a computer via USB and displayed on-screen in real time after digital low-pass filtering. The raw signal data were also stored in a file on the computer for future retrieval, processing and plotting. Data processing operations to generate the plots shown in Figure 2–Figure 6 involved simple operations in Mathematica (Wolfram Inc.) including: (i) low-pass Hamming or moving-average filters to suppress 50 Hz mains-electricity noise, which was the dominant noise source of the magnetometer, and (ii) Fourier transformation, as in Figure 5. A record of these operations is contained in the Supporting Information.
Simulations: All simulations were carried out using the SpinDynamica packages for Mathematica.[67] All simulations involved three spins, two protons and a carbon-13, and the interactions considered were the Zeeman interaction between the spins, the applied magnetic fields (in , , or ) and the spin-spin -couplings (untruncated couplings in all cases). Chemical shifts were omitted because at such low fields these terms are negligible. Relaxation was included phenomenologically for the simulations in Figure 2c and Figure 4, as a constant rate of signal decay over time, independent of field or the type of spin order.
The density operator of the fumarate/maleate systems immediately after the PHIP process was assumed to be pure nuclear-spin singlet order (of the two protons and ) secularized under the -coupling Hamiltonian plus a 50 T background field to approximate Earth’s field. The -couplings used for the simulations are shown in Figure 2c, and these couplings were modified within the given errors (50 mHz) for each simulation to fit to the data. The field profiles that are plotted in Figure 2c, Figure 3 and Figure 4 were obtained by propagating this initial density operator under the applied field sweep(s), and evaluating the projection of the sample’s magnetization along a vector tilted 30º from the axis towards the axis, at each time point. This is given by the observable:
where the operator for the total magnetization along the axis is: . The 30º rotation corresponds to the magnetometer’s sensitive axis.
A Mathematica notebook is included in the Supporting Information containing the raw simulation code with comments.
Data, Materials, and Software Availability
Experimental data files supporting the findings of this study plus a Mathematica notebook containing processing and simulation code with comments are freely available via the OPENAIRE database, hosted on Zenodo: URL https://doi.org/10.5281/zenodo.13752103. All other data are included in the manuscript and/or supporting information.